### v1

# Christian Röver, Tim Friede

Abstract
[Summary]Meta-analytic-predictive (MAP) priors have been proposed as a generic approach to deriving informative prior distributions, where external empirical data are processed to learn about certain parameter distributions. The use of MAP priors is also closely related to shrinkage estimation (also sometimes referred to as dynamic borrowing). A potentially odd situation arises when the external data consist only of a single study. Conceptually this is not a problem, it only implies that certain prior assumptions gain in importance and need to be specified with particular care. We outline this important, not uncommon special case and demonstrate its implementation and interpretation based on the normal-normal hierarchical model. The approach is illustrated using example applications in clinical medicine.
keywords:
Random-effects meta-analysis, shrinkage estimation, MAP prior, power prior, bias allowance, dynamic borrowing.
articletype: Research articlejournal: Research Synthesis Methodsarticletype: RESEARCH ARTICLE
Highlights
What is already known:
  • shrinkage estimation may be used to effectively and robustly borrow information between related data sources
  • shrinkage estimation may alternatively be motivated via a meta-analytic-predictive (MAP) approach
What is new:
  • a MAP approach remains sensible down to the extreme case of only a single study
  • the MAP prior’s usual features are retained, in addition there are connections to power prior and bias allowance approaches
Potential impact for RSM readers outside the authors’ field:
  • MAP priors are useful for constructing empirically motivated priors based on external/historical data
  • practical application is straightforward using existing software packages
1 Introduction
When evidence from a single experiment, such as a clinical trial, is not sufficiently conclusive on its own, it may sometimes help to view the data in the context of related instances (similar experiments) in order to yield more confident conclusions. This idea is explicitly implemented in shrinkage estimation, where a hierarchical model is set up accounting for estimation uncertainly at the study level as well as for variability (and similarity) between studies 1, 2; models of this kind are commonly also used in the context of meta-analysis 3, 4. The borrowing-of-information taking place between the study of primary interest and the external data may be viewed in terms of the overarching joint model as a meta-analytic-combined (MAC) approach, or, equivalently, by formulating the meta-analytic predictive (MAP) prior that explicates the information contributed by the external data to the shrinkage estimate 5. A particular special case is given when a single (“target”) study is supported by a single (“source”) study; such situations are not uncommon, and shrinkage estimation here has proven useful 6, 7, 8. Application of a hierarchical model makes it behave dynamically in the sense that more or less information is borrowed, depending on the apparent similarity of target and source data 6. When considering this case in terms of the implied MAP prior, the “meta-analysis” involved here is based on a single study, which may appear somewhat counterintuitive at first. However, such a smallest possible meta-analysis does not pose a conceptual problem; it only implies that, due to the particular sparsity of data, prior specification within the model receives special importance, a problem which, however, is common in meta-analysis of few studies in general 9], and which analogously applies for alternative borrowing methods, such as power priors or bias allowance models [8, 10. Since others appear to have struggled with or shied away from the idea of a single-study meta-analysis 11, 12, it seems worth while investigating this special case a bit closer.
The remainder of this article is structured as follows; in Section 2 the normal-normal hierarchical model (NNHM) is introduced, the meta-analysis model which then is the basis for shrinkage estimation between a pair of studies, and for the MAP prior based on a single study. The ideas will be illustrated in two practical examples in the following Section 3. Section 3.1 discusses an application in paediatric Alport syndrome that was originally formulated in terms of a shrinkage estimation problem. Section 3.2 introduces a trial design application in cardiology, where information from a similar past study is designated for consideration in the eventual analysis via an informative MAP prior. Section 4 then closes with a brief discussion.
2 Shrinkage estimation using two studies
2.1 The normal-normal hierarchical model (NNHM)
The most common model for random-effects meta-analysis is given by the normal-normal hierarchical model (NNHM). It implements sampling error as well as between-study heterogeneity using normal distributions. The data are given in terms of $k$ estimates $y_{i}$ and associated standard errors $s_{i}$ ($i=1,\ldots,k$). Each individual study aims to quantify a parameter $\theta_{i}$, so that
$y_{i}|\theta_{i},s_{i}\;\sim\;\mathrm{Normal}(\theta_{i},s_{i}^{2})\mbox{.}$ (1)
The underlying parameters $\theta_{i}$ are not necessarily identical for all studies, instead some amount of (between-study) heterogeneity is allowed for, expressed as
$\theta_{i}|\mu,\tau\;\sim\;\mathrm{Normal}(\mu,\tau^{2})\mbox{.}$ (2)
Often the overall mean $\mu$ is the aim of the analysis, while sometimes the study-specific parameters $\theta_{i}$ are also of interest 13, 6. The heterogeneity, while important, usually remains a nuisance parameter. In the context of “shrinkage estimation” of the $\theta_{i}$, an interesting aspect is that the problem may be motivated in two ways; classically, one may think of shrinkage estimation as a joint analysis of all ($k$) estimates, which also returns estimates of any $\theta_{i}$ parameter along the way; this is also denoted as the meta-analytic-combined (MAC) approach. The problem, however, may also be factored into the evidence stemming from the $i$th study alone, as well as the information provided by the remaining ($k-1$) estimates. Shrinkage estimation then may be interpreted as the analysis of the $i$th study, based on a prior distribution that results as the predictive distribution derived from a meta-analysis of the other ($k-1$) studies; this prior is denoted as the meta-analytic-predictive (MAP) prior. Both MAC and MAP approaches are equivalend and yield identical shrinkage estimates 5.
In the following, we will focus on the special case of only two studies ($k=2$). For the shrinkage estimate ($\theta_{2}$), this implies a MAP prior that is based on a single study (i.e., the data provided through $y_{1}$ and $s_{1}$). While this may appear odd at first, the idea readily applies also in this special case, as will be demonstrated in the following. Analysis may generally be performed based on informative or uninformative priors for $\mu$, while a proper, informative prior is required for $\tau$ 4, 9. The case of only $k=2$ studies is also closely connected to the related concepts of power prior 6 or bias allowance models 14, 10.
2.2 Uniform effect prior
Priors for the overall mean parameter ($\mu$) in the NNHM may be specified as informative or as uninformative. For (more or less informative) priors, normal distributions are an obvious choice, also since these lead to analytically simple inference. Quite commonly, effect priors however are chosen as uninformative and (improper) uniform, not least due to certain analogies to frequentist meta-analysis procedures 4. In case of an (improper, non-informative) uniform prior for $\mu$, certain expressions turn out particularly simple, which is why we will focus on this particular, yet insightful, common and practically relevant case in the following. As the (improper) uniform prior constitutes the limiting case of an increasingly uninformative effect prior, the following considerations may also be viewed as relating to the limiting behaviour for increasingly uninformative priors (e.g., for normal priors when their variance approaches infinity). The uniform effect prior leads to a normal conditional posterior for the overall mean $\mu$, with moments given by
$\mathrm{E}[\mu|y_{1},s_{1},\tau]\;=\;y_{1}\quad\mbox{and}\quad\mathrm{Var}(\mu|y_{1},s_{1},\tau)\;=\;s_{1}^{2}+\tau^{2}\mbox{,}$ (3)
and a marginal heterogeneity likelihood that is constant (independent of $\tau$), so that the heterogeneity’s posterior equals its prior 4. This seems reasonable, since (as long as the effect prior is uniform) a single observation $y_{1}$ does not provide information on the heterogeneity $\tau$.
2.3 The MAP prior for the effect in a new study
The MAP prior results as the posterior predictive distribution for a “new” (second) study’s effect (here: $\theta_{2}$) given the data from the first study ($y_{1}$, $s_{1}$). In the NNHM framework, the conditional predictive distribution again is normal with mean
$\mathrm{E}[\theta_{2}|y_{1},s_{1},\tau]\;=\;\mathrm{E}[\mu|y_{1},s_{1},\tau]\;=\;y_{1}$ (4)
and variance
$\mathrm{Var}(\theta_{2}|y_{1},s_{1},\tau)\;=\;\mathrm{Var}(\mu|y_{1},s_{1},\tau)+\tau^{2}\;=\;s_{1}^{2}+2\tau^{2}$ (5)
(see also (3) and (2)). These expressions make sense in the present context: We know $\theta_{1}$ with accuracy given by the standard error $s_{1}$, and we know that the difference between $\theta_{1}$ and $\theta_{2}$ is normally distributed with variance $2\tau^{2}$, so that the (conditional) variance expression results as a corresponding sum 6. As pointed out above, information on heterogeneity ($\tau$) is based on the prior only; the heterogeneity in this context generally requires a proper, informative prior (since $k<3$) 4, 9.
The eventual (marginal) predictive distribution (marginalized over the distribution of $\tau$) hence results as a normal scale mixture 15, 16 with fixed mean (4) and with variance as given in (5), where $\tau$ is distributed according to the specified prior. In particular, one may think of the MAP prior as the first study’s point estimate $p(\theta_{1}|y_{1},s_{1})$ convolved with the prior predictive distribution $p(\theta_{2}|\theta_{1},\tau)$ and then marginalized over $\tau$. The MAP prior is symmetric around $y_{1}$, and its (marginal) variance results from (5) as
$\mathrm{Var}(\theta_{2}|y_{1},s_{1})\;=\;\mathrm{E}[s_{1}^{2}+2\tau^{2}]\;=\;s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]\mbox{,}$ (6)
where the expectation $\mathrm{E}[\tau^{2}]$ depends on the assigned heterogeneity prior. For a range of common prior specifications, this expectation may be derived analytically; Table 3 in Appendix A.2 lists some popular cases. From this expression, one can see that the relative magnitudes of $s_{1}^{2}$ and $\mathrm{E}[\tau^{2}]$ determine whether the resulting MAP prior’s variance is dominated by estimation uncertainty (regarding $\theta_{1}$) or anticipated heterogeneity ($\tau$). These two variance components may in fact also be considered to reflect so-called “type A” and “type B” uncertainties relating to measurement uncertainty and background knowledge, respectively, which together sum up to form the combined standard uncertainty $u_{c}$ (the square root of (6)) 17, 18.
Since the MAP prior results as a normal scale mixture, it is generally heavier-tailed than a normal distribution, which has implications for the resulting operating characteristics. A heavy-tailed (MAP-) prior means that in combination with a (“shorter-tailed”) normal likelihood, the likelihood will dominate in case of a prior-data conflict 19. Such robustness properties have in fact been noted and demonstrated in the meta-analysis context, as these lead to a dynamic borrowing behaviour 6. The MAP priors’ tail behaviour will also be illustrated for some examples below (Figures 4 and 5).
Another way to quantify the precision of a prior distribution is by translating it into a number of observations (effective sample size, ESS) that would in a certain sense convey an equivalent amount of information; in the following, we will quote ESSs based on the expected local-information ratio ($\mathrm{ESS}_{\mathrm{ELIR}}$) 20.
2.4 The bias allowance model connection
In the 2-study case, there is a one-to-one correspondence between the NNHM and a simple bias allowance model 10; instead of the NNHM assumption (2) in combination with a uniform effect prior and heterogeneity prior $p_{\star}(\tau)$ as in Section 2.1, one may specify
$\theta_{2}|\alpha,\beta$ $=$ $\alpha$ (7)
$\theta_{1}|\alpha,\beta$ $\sim$ $\mathrm{Normal}(\alpha,\,\beta^{2})$ (8)
with prior $p(\beta)=\frac{1}{\sqrt{2}}\,p_{\star}\bigl{(}\frac{\beta}{\sqrt{2}}\bigr{)}$ for the standard deviation $\beta$ 6. This “reference model” is different in that one estimate (the reference, or “target” $y_{2}$) directly relates to $\alpha$, while the other one (the “source” $y_{1}$) is associated with an additional offset to account for potential bias. The shrinkage estimates of $\theta_{i}$, however, can be shown to be identical in both models as long as a uniform prior for the effect ($\mu$) is used 6. The reference model may be considered a variation of Pocock’s bias model or, more generally, a bias allowance model 21, 14, 10. The $\beta$ parameter, which only differs from $\tau$ by a scaling factor of $\sqrt{2}$, may also help motivating a heterogeneity prior, as it directly relates to the expected difference between $\theta_{2}$ and $\theta_{1}$, without reference to a common overall mean $\mu$.
2.5 The power prior connection
There is also is a connection to a so-called power prior, which has been proposed as an approach for deliberate down-weighting of prior information. It is intended for a prior distribution that itself results as a posterior, and the power prior results from applying an exponent $a_{0}$ (with $0\leq a_{0}\leq 1$) to its likelihood contribution 22, 21. When conditioning on a fixed $\tau$ value, the (conditional) MAP prior is normal with moments given in (4) and (5); in particular, note that $\tau^{2}$ acts additively on the “plain” variance ($s_{1}^{2}$). In this context, a power prior with fixed exponent $a_{0}$ on the other hand would correspond to a $\mathrm{Normal}\Bigl{(}y_{1},\,\frac{s_{1}^{2}}{a_{0}}\Bigr{)}$ distribution, where the (inverse) exponent acts multiplicatively on the variance. Both MAP and power prior then are identical if $a_{0}=\bigl{(}2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr{)}^{-1}$ 6, 23. While a prior probability distribution for $\tau$ is readily motivated with reference to the effect scale of $y_{i}$ and $\theta_{i}$ 9, specification of a fixed $\alpha_{0}$ value remains tricky, and a prior distribution may be even harder to motivate. On the other hand, through the above functional correspondence, any prior for the heterogeneity $\tau$ immediately implies a corresponding distribution for the exponent $a_{0}$; an example is illustrated in Appendix A.3.
3 Two practical applications
3.1 Pediatric Alport example
3.1.1 Application
Gross et al. (2020) 24 performed a randomized controlled trial (RCT) in Alport syndrome to investigate the effects of ramipril, an angiotensin-converting enzyme inhibitor (ACEi). Recruitment of participants to the RCT was hampered by the rare and pediatric nature of the disease, and so the analysis of the RCT had been planned with the inclusion of observational data from an open-label arm and a natural disease cohort 25. Time to disease progression was a co-primary endpoint, and from the observational data, a hazard ratio (HR) of 0.53 [0.22, 1.29] was estimated based on 70 patients. Only 20 patients entered into the RCT, and a HR of 0.51 [0.12, 2.20] was estimated. The data are also summarized in Table 1.
Table 1: Alport example data from Gross et al. (2020) 24
log-HR
$i$ estimate $n_{i}$ hazard ratio (HR) $y_{i}$ $s_{i}$
1 observational 70 0.53 [0.22, 1.29] $-0.635$ $0.451$
2 RCT 20 0.51 [0.12, 2.20] $-0.673$ $0.742$
The analysis was then performed by jointly considering both log-HR estimates in a NNHM, anticipating a reasonable amount of heterogeneity between them (expressed though a $\mathrm{half\mbox{-}Normal}(0.5)$ prior for $\tau$), and deriving a shrinkage estimate for the RCT effect ($\theta_{2}$). The resulting estimate then was substantially more precise than if the RCT data were considered in isolation; the HR was estimated at 0.52 [0.19, 1.39] 24.
To make the flow of information transparent, we may now derive the corresponding MAP prior reflecting the information contributed by the observational data.
Figure 1: Illustration of MAP-prior, likelihood and (shrinkage-) posterior for the Alport example discussed in Section 3.1 24. The horizontal lines at the bottom indicate point estimates and corresponding 95% intervals
Figure 1 illustrates MAP-prior, likelihood and posterior (shrinkage estimate) for the Alport example. The MAP prior here has a mean of $y_{i}=-0.63$ and a variance of $s_{1}+2\,\mathrm{E}[\tau^{2}]=0.45^{2}+2\times 0.5^{2}=0.84^{2}$. While the observational sample size was 70 patients, the MAP prior’s effective sample size ($\mathrm{ESS}_{\mathrm{ELIR}}$) is at only 26 patients (i.e., 37% of originally 70 actual patients) 20. The eventual shrinkage interval is only 67% as wide as the original, implying a substantial “effective gain in sample size” 6; such a precision increase would otherwise have required more than doubling the sample size (by the addition of 24 extra patients).
3.1.2 Variations of the MAP prior
The heterogeneity ($\tau$) prior was specified as a $\mathrm{half\mbox{-}Normal}(0.5)$ distribution, which is a reasonably conservative choice for endpoints such as HRs, as it covers “reasonable” and up to “fairly high” levels of heterogeneity ($\tau\leq 1$) and leaves a small prior probability for “fairly extreme” amounts ($\tau>1$) 26, 9. Since conclusions heavily depend on the heterogeneity prior settings, it may however be interesting to investigate the effects of alternative specifications; in particular, we will consider different prior scales and different distribution families.
Assuming that $s_{1}=0.451$ (as in the present example, see Table 1), we can illustrate the resulting MAP prior when varying the heterogeneity prior scale.
Figure 2: Illustration of the resulting MAP-prior for varying heterogeneity prior scales. The dashed line indicates the likelihood of the observational data alone for comparison
Figure 2 shows the likelihood of the observational estimate along with the corresponding MAP priors for half-normal heterogeneity priors with scales 0.25, 0.50 and 1.00. Increasing the heterogeneity prior scale yields a MAP prior that becomes increasingly wider than the plain likelihood alone. In the present case, the effect scale was a logarithmic HR, and on the logarithmic scale, the original MAP prior (based on a $\mathrm{half\mbox{-}Normal}(0.5)$ heterogeneity prior) covers a range of $y_{1}\pm 1.72$ with 95% probability. On the exponentiated scale (see also the top axis in Figure 2), a difference of 1.72 in log-HR would correspond to a 5.6-fold larger HR. If one switched to a $\mathrm{half\mbox{-}Normal}(0.25)$ or $\mathrm{half\mbox{-}Normal}(1.0)$ prior instead, the range would change to $y_{1}\pm 1.13$ or $y_{1}\pm 3.18$ instead, corresponding to multiplicative factors of 3.1 or 24.0, respectively.
Half-normal distributions are a common and obvious choice as heterogeneity priors, possible reasons may be familiarity and availability, as well as a “flat” shape near the origin and a rather short tail 9. Variations of the distribution family commonly do not alter conclusions dramatically as long as they cover a similar range, as manifested e.g. in a common prior median.
MAP priors corresponding to alternative specifications to a $\mathrm{half\mbox{-}Normal}(0.5)$ for the heterogeneity are illustrated in Figure 3. A range of distribution families are used, with their scale parameters specified such that all correspond to a common prior median for $\tau$ (of $0.34$). These different heterogeneity prior families are also shown in Appendix A.1.
Figure 3: Illustration of MAP-prior’s dependence on heterogeneity prior distribution family. The different heterogeneity priors shown here all share the same prior median
The resulting MAP prior densities themselves are hard to distinguish. Differences are more noticeable when focusing on the tail behaviour, e.g., considering cumulative distribution functions (as shown in Figure 4) or logarithmic densities (shown in Figure 5).
Figure 4: The MAP-priors’ cumulative distribution functions corresponding to the densities shown in Figure 5
Figure 5: The MAP-priors’ densities on a logarithmic scale (see also Figure 5). Note that the likelihood for the observational data alone follows a parabola shape here, while the corresponding MAP priors are clearly much heavier-tailed
One can see that heavier-tailed heterogeneity priors also yield correspondingly heavier-tailed MAP-priors. The heterogeneity and the corresponding MAP priors’ properties are also summarized and compared in Table 2 in terms of prior quantiles and effective sample sizes ($\mathrm{ESS}_{\mathrm{ELIR}}$).
Table 2: Summaries of MAP priors resulting from several settings for the heterogeneity ($\tau$) prior. The half-normal(0.5) prior is contrasted with half-normal priors of differing scale, as well as with priors of differing distributional families, but with matching prior medians. Note that in the context of the present example, the MAP prior’s domain corresponds to logarithmic hazard ratios (log-HRs). Quantiles are centered at $y_{1}$
$\tau$ prior standard quantile ($\theta_{2}-y_{1}$)
family scale median $\mathrm{ESS}_{\mathrm{ELIR}}$ deviation 95.0% 97.5% 99.5%
half-normal 0.50 0.34 26.6 0.84 1.32 1.72 2.72
half-normal 0.25 0.17 45.7 0.57 0.93 1.13 1.62
half-normal 1.00 0.67 12.8 1.48 2.35 3.18 5.19
half-$t_{\nu=4}$ 0.46 0.34 25.3 1.02 1.45 1.98 3.58
half-Cauchy 0.34 0.34 23.4 2.45 4.85 24.02
half-logistic 0.31 0.34 25.8 0.91 1.39 1.85 3.09
exponential 0.49 0.34 24.5 1.07 1.56 2.19 3.96
Lomax($\alpha\!=\!6$) 2.75 0.34 24.0 1.31 1.70 2.50 5.05
Lomax($\alpha\!=\!1$) 0.34 0.34 23.1 3.29 7.05 37.17
3.2 Heart failure example
The Spirit-HF trial has been designed in order to test the efficacy of spironolactone in patients with heart failure (HF) 27. Spironolactone is expected to reduce cardiovascular mortality as well as hospitalizations due to heart failure, and had previously been investigated in the Topcat trial 28. Both studies refer to the composite of (recurrent) HF hospitalization and cardiovascular death as the primary endpoint to evaluate treatment efficacy. Despite a sizeable sample size of 3445 patients and a mean follow-up duration of more than $3$ years, the Topcat trial failed to demonstrate statistical significance; the estimated hazard ratio (HR) was at 0.89 (0.77, 1.04) ($p=0.14$) 28.
The analysis of the new Spirit-HF trial meanwhile is being planned, and may take into consideration the evidence already generated in the Topcat trial. One idea may be to derive a shrinkage estimate, anticipating some between-study heterogeneity, and dynamically borrowing information from the earlier study based on the corresponding MAP prior 6, 7. For the between-study heterogeneity $\tau$, use of a $\mathrm{half\mbox{-}Normal}(0.25)$ prior may be appropriate. The heterogeneity prior may be motivated referring to anticipated levels of heterogeneity based on general considerations 9 or using empirical evidence, in particular in view of the similar study designs and the effect measure being a log-HR 29.
In the present case, analysis is based on the logarithmic HR; the HR estimated in the Topcat trial corresponds to a log-HR of $-0.117$ with a standard error of $0.077$. The corresponding Topcat likelihood along with the resulting MAP prior is illustrated in Figure 6.
Figure 6: Illustration of likelihood and corresponding MAP-prior for the heart failure example, using a $\mathrm{half\mbox{-}Normal}(0.25)$ prior for $\tau$. The horizontal line at the bottom indicates the 95% prediction interval
The variance (squared standard error) of the Topcat study’s estimate was $s_{1}^{2}=0.077^{2}$ while for the assumed heterogeneity prior the expected heterogeneity variance is $\mathrm{E}[\tau^{2}]=0.25^{2}$ (see also Table 3), so that the resulting MAP prior’s variance (6) is $s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]=0.362^{2}$, and the majority of the variance is due to epistemic uncertainty relating to the anticipated similarity of the Topcat and Spirit-HF estimates. The 95% prediction interval for the MAP prior is centered at the Topcat log-HR estimate and ranges from $-0.899$ to $+0.665$, corresponding to HRs in the range [$0.407$, $1.945$]. According to the MAP prior, the probability of a beneficial treatment effect (a log-HR below zero) is 71%. The MAP prior has an ESSELIR of 399, i.e, only 12% of the 3445 actual Topcat patients, and 31% of the estimated enrolment of 1300 Spirit-HF patients 27.
4 Discussion
Despite the seemingly odd notion of a meta-analysis of a single study, the use of MAP priors remains completely consistent down to the extreme case of only one data point. The “usual” toolbox remains available, including common prior specifications 9, computation of effective sample sizes (ESS) 20 robustification 5, as well as common meta-analysis software (e.g., the bayesmeta or RBesT R packages) 4, 30 for practical implementation. In addition, for $k=1$, there are connections to bias allowance and power prior models (see Section 2) that may help motivating such an approach. MAP priors based on few estimates are generally rather heavy-tailed, which will ensure robust operating characteristics 19, 6, 7. For few data points in general, and in particular for only a single data point, the prior specification for the heterogeneity parameter $\tau$ gains in importance and needs to be particularly well-founded and convincing 9.
Empirical MAP priors may then be utilized in different ways, either to simply motivate a reasonable sample size (or other design aspects), or to implement explicit borrowing of historical information 31, 32. When MAP priors are used to also inform the analysis, it is important to approach the evaluation of operating characteristics from a sensible angle 33, 34, 35.
Van Zwet et al. (2024) argue that also the analysis of a single study should account for heterogeneity of the treatment effect across studies. Therefore, they propose to consider analyses of individual studies also within an overarching NNHM framework similar to our approach presented here; using informative, empirically motivated priors for both $\mu$ and $\tau$, inference may then be focused on the overall mean effect ($\mu$) rather than the study-specific $\theta_{1}$ even in the analysis of only a single study 36.
Appendix A Appendix
A.1 Heterogeneity priors
Figure 7 shows several prior densities discussed in Section 3.1.2 that are all scaled to a common median (of 0.34, the median of a HN(0.5) distribution).
Figure 7: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their probability density functions. All priors are scaled such that they have a common median of 0.34 (the median of a half-normal(0.5) prior; dashed line)
Figure 8: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their cumulative distribution functions (CDFs). All priors are scaled such that they have a common median of 0.34 (dashed line)
A.2 Variance of the MAP prior
The prior predictive distribution $p(\theta_{2}|y_{1},s_{1})$ is a normal scale mixture with (constant) mean $y_{1}$ and (conditional) variance $s_{1}^{2}+2\tau^{2}$, where $\tau$ is distributed according to the specified heterogeneity prior. The mixture distribution’s marginal variance results as $\mathrm{Var}(\theta_{2}|y_{1},s_{1})=s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]$ and depends on the (prior) expectation of the squared heterogeneity $\mathrm{E}[\tau^{2}]$ (see (6)).
Table 3: Expected values of $\tau^{2}$ based on various common (prior) distributions for $\tau$, depending on their scale parameter $s$. An asterisk ($\ast$) indicates that there is no simple analytical expression. The half-Cauchy prior would be an additional option, but does not have a finite expectation (it is in fact also a special case of the half-Student-$t$ prior, with $\nu=1$ degree of freedom)
standard deviation $\tau$ variance $\tau^{2}$
distribution median 95% quantile expectation expectation
$\tau\sim\mathrm{half\mbox{-}Normal}(s)$ $\approx 0.674s$ $\approx 1.96s$ $\sqrt{\frac{2}{\pi}}\,s\approx 0.798s$ $s^{2}$
$\tau\sim\mathrm{half}\mbox{-}\mathrm{Student}\mbox{-}t_{\nu}(s)$ast $\ast$ $2\sqrt{\frac{\nu}{\pi}}\frac{\Gamma(\frac{\nu+1}{2})}{\Gamma(\frac{\nu}{2})(\nu-1)}s$ $(\nu>1)$ $\frac{\nu}{\nu-2}s^{2}$ $(\nu>2)$
$\tau\sim\mathrm{half}\mbox{-}\mathrm{logistic}(s)$ $\log(3)\,s\approx 1.10s$ $\approx 3.66s$ $\log(4)\,s\approx 1.39s$ $\frac{\pi^{2}}{3}s^{2}\approx 3.29s^{2}$
$\tau\sim\mathrm{Exp}(s)$ $\log(2)\,s\approx 0.693s$ $\approx 3.00s$ $s$ $2s^{2}$
$\tau\sim\mathrm{Lomax}(\alpha,s)$ $(2^{\frac{1}{\alpha}}-1)\,s$ $(20^{\frac{1}{\alpha}}-1)\,s$ $\frac{1}{\alpha-1}s$ $(\alpha>1)$ $2\frac{\Gamma(\alpha-2)}{\Gamma(\alpha)}s^{2}$ $(\alpha>2)$
$\tau\sim\mathrm{Uniform}[0,s]$ $\frac{s}{2}$ $0.95s$ $\frac{s}{2}$ $\frac{1}{3}s^{2}$
Table 3 below summarizes expectations for $\tau^{2}$ for a range of common heterogeneity priors. Note also the related Table B1 in the appendix of Röver et al. (2021) 9 giving some additional details for the prior distributions shown here.
A.3 Power prior exponent’s distribution
For any fixed heterogeneity value $\tau$, the (conditional) MAP prior is equivalent to a power prior with exponent $a_{0}=\bigl{(}2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr{)}^{-1}$ (see also Section 2.5) 6. Through this functional relationship, any prior density $p_{\star}(\tau)$ for the heterogeneity implies a corresponding prior for the exponent $a_{0}$ with probability density function
$p(a_{0})\;=\;\frac{s_{1}}{2\sqrt{2}}\frac{\sqrt{\frac{a_{0}}{1-a_{0}}}}{a^{2}}\;p_{\star}\biggl{(}s_{1}\sqrt{\frac{1-a_{0}}{2\,a_{0}}}\biggr{)}\mbox{.}$ (9)
Figure 9 illustrates such densities for an example value of $s_{1}=0.451$ (as in the Alport example from Section 3.1).
Figure 9: Illustration of prior distributions for the power prior exponent $a_{0}$ corresponding to certain prior distributions assumed for the heterogeneity $\tau$ (and $s_{1}=0.451$)
The prior densities for $a_{0}$ are shown for half-normal priors with scales 0.25, 0.50 and 1.00 (corresponding to the cases also illustrated in Figure 2).
A value of $a_{0}\!=\!1$ for the exponent corresponds to full borrowing, while smaller values imply increasing degrees of discounting of prior information. While the $\mathrm{half\mbox{-}Normal}(0.25)$ prior places a substantial share of prior probability near $a_{0}\!=\!1$, larger prior scale parameters correspond to less a-priori expected borrowing, eventually resulting in bimodal priors for $a_{0}$.
A.4 Example R code
A.4.1 Alport example
# data from:
#   Gross et al. (2020)
#   https://doi.org/10.1016/j.kint.2019.12.015
#   Figure 2(b)
GrossEtAl2020 <- cbind.data.frame("data"     = c("observational", "RCT"),
                                  "patients" = c(70, 20),
                                  "events"   = c(29, 8),
                                  "hr"       = c(0.53, 0.51),
                                  "lower"    = c(0.22, 0.12),
                                  "upper"    = c(1.29, 2.20))
GrossEtAl2020\$loghr <- log(GrossEtAl2020\$hr)
GrossEtAl2020\$se    <- (log(GrossEtAl2020\$upper)-log(GrossEtAl2020\$lower))/(2*qnorm(0.975))
# show data (see Table 1):
GrossEtAl2020
# convert effect measures to "escalc" object:
library("metafor")
es <- escalc(measure="GEN",
             yi=loghr, sei=se, ni=patients,
             slab=data, data=GrossEtAl2020)
# perform joint meta-analysis of both estimates (MAC approach):
library("bayesmeta")
bm <- bayesmeta(es,
                tau.prior=function(t){dhalfnormal(t, scale=0.50)})
# show data as well as overall and shrinkage estimates:
forestplot(bm, predict=FALSE,
           xlab="log-HR",
           txt_gp = fpTxtGp(ticks = gpar(cex=1), xlab = gpar(cex=1)))
# check out shrinkage estimate(s):
bm\$theta  # (log-HRs)
exp(bm\$theta[c("median","95% lower", "95% upper"),"RCT"])  # (HR)
# meta analysis of observational data only (MAP approach):
map <- bayesmeta(es[1,],
                 tau.prior=function(t){dhalfnormal(t, scale=0.50)})
# check out predictive distribution (MAP prior):
map\$summary[,"theta"]  # (log-HR)
# illustrate prior, likelihood, posterior / shrinkage estimate
# (see Figure 1):
colvec <- c("map"    = "red3",
            "likeli" = "blue",
            "shrink" = "green3")
x <- seq(-4, 2.5, length=200)
plot(c(-3.5, 2.0), c(0.0, 0.85), type="n",
     main="Alport example (Gross et al., 2020)",
     xlab="log-HR", ylab="")
# MAP prior (predictive density):
lines(x, map\$dposterior(theta=x, predict=TRUE), type="l", col=colvec[1])
# RCT likelihood:
lines(x, dnorm(x, mean=es\$yi[2], sd=sqrt(es\$vi[2])), type="l", col=colvec[2])
# posterior (=shrinkage estimate):
lines(x, bm\$dposterior(theta=x, individual="RCT"), type="l", col=colvec[3])
abline(h=0, col="grey20")
lines(c(0,0), c(0, 1), col="grey20", lty=2)
legend("topleft", c("MAP prior (observational)", "likelihood (RCT)",
                    "shrinkage (RCT)"),
       col=colvec, pch=15)
A.4.2 Heart failure example
# TOPCAT study (Pitt et al., 2024) quotes a HR of:
# 0.89 (0.77, 1.04)
# log-HR:
log(0.89)  # = -0.117
# standard error:
(log(1.04) - log(0.77)) / (2*qnorm(0.975))  # = 0.077
# derive MAP prior:
library("bayesmeta")
bm <- bayesmeta(y=-0.117, sigma=0.077, label="Topcat",
                tau.prior=function(t){dhalfnormal(t, scale=0.25)})
# (MAP-) prior predictive interval:
round(rbind("log-HR"=bm\$summary[c(2,5:6), "theta"],
            "HR"=exp(bm\$summary[c(2,5:6), "theta"])), 3)
# prior probability of HR<0:
bm\$pposterior(theta=0, predict=TRUE)
# illustrate TOPCAT likelihood
# and associated MAP prior (see Figure 6):
x <- seq(-1.8, 1.8, length=200)
plot(c(-1.3, 1.3), c(0.0, 5.18), type="n",
     xlab="log-HR", ylab="")
# TOPCAT likelihood:
lines(x, dnorm(x, mean=-0.117, sd=0.077), lty=2)
# MAP prior:
lines(x, bm\$dposterior(theta=x, predict=TRUE), col="red3", lwd=1.5)
# axes, legend:
abline(h=0, v=0, col="grey60")
legend("topright", lty=c(2,1), col=c("black", "red3"), lwd=c(1,1.5),
       c("TOPCAT likelihood", "MAP prior"))
# compute TOPCAT’s unit information standard deviation (UISD):
# (standard error was 0.077, based on 3445 patients)
uisd(n=3445, sigma=0.077)  # = 4.5
# compute MAP prior’s effective sample size (ESS_ELIR):
ess(bm, uisd=4.5)  # = 399 patients
Funding Statement
Support from the German Centre for Cardiovascular Research (Deutsches Zentrum für Herz-Kreislauf-Forschung e.V., DZHK) is gratefully acknowledged (grant number 81Z0300108).
Competing Interests
The authors have declared no conflict of interest.
Data Availability Statement
The data and code that supports the findings of this study are available in the supplemental material of this article.
Author Contributions
Conceptualization: T.F.; C.R. Methodology: C.R. Writing original draft: C.R. All authors approved the final submitted draft.